The weights (in kg) of Grade 11 ICT follow a normal distribution and has a mean of 52 and standard deviation of 6. How many students have weights greater than 55?


1
Expert's answer
2022-05-30T15:01:03-0400
"P(X>55)=1-P(Z\le\frac{55-52}{6})""=1-P(Z\le0.5)\approx0.3085"

30.85 %